OpenStudy (anonymous):

r-6sintheta=0 find equation that is in x and y form

OpenStudy (freckles):

\[r=6 \sin(\theta) \\ \text{ hint multiply r on both sides }\]

OpenStudy (freckles):

then use any of the following that help to write the equation in cartesian form: \[r^2=x^2+y^2 \\ r \sin(\theta)=y \\ r \cos(\theta) =x \\ \tan(\theta)=\frac{y}{x}\]

OpenStudy (anonymous):

so then r^2=6rsintheta

OpenStudy (anonymous):

r^2=6y right?

OpenStudy (freckles):

and r^2=x^2+y^2 :)

OpenStudy (anonymous):

oh yeah! ok then you complete the square right?

OpenStudy (freckles):

if you need to write it in the standard form for a circle yes then you would need to complete the square

OpenStudy (freckles):

and sorry my notifications aren't working properly